OpenStudy (anonymous):

solve for x 3(-2x-1)=2(x+5)

OpenStudy (anonymous):

You must isolate the variable.

OpenStudy (anonymous):

-6x+3=2x+10 What is multiplied +6x +6x Add 6x to both sides __________________ 3 = 8x+10 Subtract 10 to isolate the variable X -10 -10 __________________ -7/8 = 8x/8 Divide by 8 x = -7/8

OpenStudy (anonymous):

thank u